OpenStudy (anonymous):

1) Find dy/dx ]x=4 if y=u^2-2u^5 and u=x-sqrt x. 2) Find dy/dx]t=1 if y=sqrt(1+r^2) and r= (t+1)/(2t+1)

OpenStudy (dumbcow):

this is chain rule \[\frac{dy}{dx} = \frac{dy}{du}*\frac{du}{dx}\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

do i sub in the x value into the u equation?

OpenStudy (anonymous):

or do the chain rule first?

OpenStudy (anonymous):

@freckles Can you help me please?

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

` y=u^2-2u^5 and u=x-sqrt x.` \[\frac{dy}{dx} = \frac{dy}{du}*\frac{du}{dx}\] first find \(\dfrac{dy}{du}\) and \(\dfrac{du}{dx}\)

ganeshie8 (ganeshie8):

\(y = u^2 - 2u^5\) \(\dfrac{dy}{du} = ?\)

OpenStudy (anonymous):

(2u-10u^4)

ganeshie8 (ganeshie8):

\(u = x-\sqrt{x}\) \(\dfrac{du}{dx} = ?\)

OpenStudy (anonymous):

(1-1/2x^-1/2)

ganeshie8 (ganeshie8):

so by chain rule we have \[\begin{align}\frac{dy}{dx} &= \frac{dy}{du}*\frac{du}{dx}\\~\\ &= \left(2u-10u^4\right)*\left(1-\frac{1}{2}x^{-1/2}\right) \end{align}\]

ganeshie8 (ganeshie8):

you want to find the derivative at x = 4 so plugin x = 4 above

OpenStudy (anonymous):

okay, so the answer would be -116?

ganeshie8 (ganeshie8):

try again

ganeshie8 (ganeshie8):

I'm getting -117

ganeshie8 (ganeshie8):

just double check once..

OpenStudy (anonymous):

oh i forgot to put brackets, i got the -117 too.

OpenStudy (freckles):

me too! :)

OpenStudy (freckles):

does 2 have type-o?

OpenStudy (freckles):

dy/dt]t=1 ?

OpenStudy (anonymous):

yes, that' what i mean. thanks

OpenStudy (anonymous):

So first we do the chain rule?

OpenStudy (freckles):

yah

OpenStudy (freckles):

\[\frac{dy}{dt}=\frac{dr}{dt} \cdot \frac{dy}{dr}\]

OpenStudy (anonymous):

is this correct? y' = 1/2 (1+r^2)(2r) ((-1(2t+1)(2))

ganeshie8 (ganeshie8):

\(y=\sqrt{1+r^2}\) \(\dfrac{dy}{dr} = ?\)

OpenStudy (anonymous):

oops... is it 1/2 (1+r^2)^-1/2 (2r)

ganeshie8 (ganeshie8):

Yes! \( r= \dfrac{t+1}{2t+1}\) \(\dfrac{dr}{dt} = ?\) (you need to use quotient rule)

ganeshie8 (ganeshie8):

heard of quotient rule before ? if not, you may use product rule also by writing 1/(2t+1) as (2t+1)^(-1)

OpenStudy (anonymous):

I know the quotient rule, is it faster to use that?

ganeshie8 (ganeshie8):

both are same but i would use qyotient rule here..

ganeshie8 (ganeshie8):

because i see a quotient here..

ganeshie8 (ganeshie8):

\[\large \left(\dfrac{u}{\color{blue}{v}}\right)' = \dfrac{\color{blue}{v}u' - u\color{blue}{v}'}{\color{blue}{v}^2}\]

OpenStudy (anonymous):

is this correct? \[\frac{ t+1-(t+1)(-2(2t+1)) }{ (2t+1)^2 }\]

ganeshie8 (ganeshie8):

Nope. numerator has a mistake you need to start numerator with the term in denominator : \[\large \left(\dfrac{u}{\color{blue}{v}}\right)' = \dfrac{\color{blue}{\bbox[4px,border:2px solid red]{v}}u' - u\color{blue}{v}'}{\color{blue}{\bbox[4px,border:2px solid red]{v}}^2}\] \(\)

ganeshie8 (ganeshie8):

\[\begin{align} \left(\dfrac{t+1}{\color{blue}{2t+1}}\right)' &= \dfrac{\color{blue}{\bbox[4px,border:2px solid red]{(2t+1)}}(t+1)' - (t+1)\color{blue}{(2t+1)}'}{\color{blue}{\bbox[4px,border:2px solid red]{(2t+1)}}^2}\\~\\~\\~\\~\\ &= \dfrac{\color{blue}{(2t+1)}(1) - (t+1)\color{blue}{(2)}}{\color{blue}{(2t+1)}^2}\\~\\~\\~\\ &= \dfrac{\color{blue}{2t+1} - 2t-2}{\color{blue}{(2t+1)}^2}~\\~\\~\\~\\ &= \dfrac{-1}{\color{blue}{(2t+1)}^2}~\\~\\~\\~\\ \end{align}\]

ganeshie8 (ganeshie8):

so we have below : \[\dfrac{dy}{dr} =\frac{1}{2} (1+r^2)^{-1/2} (2r) \] \[\dfrac{dr}{dt} =\frac{-1}{(2t+1)^2} \] multiply both and obtain \(\dfrac{dy}{dt}\)

ganeshie8 (ganeshie8):

\[\dfrac{dy}{dt} = \dfrac{dy}{dr}*\dfrac{dr}{dt}\]

OpenStudy (anonymous):

\[\frac{ -2r }{ 2\sqrt{1+r^2}(2t+1)^2 }\]

OpenStudy (anonymous):

@ganeshie8